# On the Role of Attention Heads in Large Language Model Safety

### v1

@@ 4.3 How Does Safety Heads Affect Safety?
Our Method Localizes Safety Parameters at a Finer Granularity. Previous research on interpretability (Zou et al., 2023a; Xu et al., 2024c), such as ActSVD (Wei et al., 2024b), Generation-Time Activation Contrasting (GTAC) & Dynamic Activation Patching (DAP) (Chen et al., 2024) and Layer-Specific Pruning (LSP) (Zhao et al., 2024b), has identified safety-related parameters or representations. However, our method offers a more precise localization, as detailed in Table 1. We significantly narrow down the focus from parameters constituting over 5% to mere 0.006%, improving attribution precision under similar ASR by three orders of magnitude compared to existed methods.
@@ 5.1 Different Impact between Attention Weight and Attention Output
Table 3: The impact of the number of ablated safety attention heads on ASR. Upper. Results of attributing safety heads at the dataset level using generalized Ships; Bottom. Results of attributing attributing specific harmful queries using Ships.
@@ 5.2 Pre-training is Important For LLM Safety
To explore the association between safety within attention heads and the pre-training phase, we conduct an experiment where we load the attention parameters from the base model while keeping the rest of the parameters from the aligned model. We then evaluate the safety of this ‘concatenated’ model and discover that it retains safety capability very close to that of the aligned model, as shown in Figure 6(a). This observation further supports the notion that the safety effect of the attention mechanism is primarily derived from the pre-training phase. Specifically, reverting parameters to their pre-alignment state does not significantly diminish safety capability, whereas ablating a safety head does.
@@ 5.3 Helpful-Harmless Trade-off
The neurons in LLMs exhibit superposition and polysemanticity (Templeton, 2024), meaning they are often activated by multiple forms of knowledge and capabilities. Therefore, we evaluate the impact of safety heads ablation on helpfulness. We use lm-eval (Gao et al., 2024) to assess model performance on zero-shot tasks after ablating safety heads. As shown in Figure 6(b), we find that safety head ablation significantly degrades the safety capability while causing little helpfulness compromise. Based on this, we argue that the safety head is indeed primarily responsible for safety.
@@ A.1 Undifferentiated Attention
Finally, the output $h_{i}$ of the attention head degenerates to $\frac{1}{N}W_{v}^{i}$, which holds exactly when $\epsilon=0$.
@@ A.2 Modifying the Value Matrix Reduces the Contribution
In Figure 8, we visualize the attention weights after applying the mean operation to Query ($W_{q}$), Key ($W_{k}$), and Value ($W_{v}$), as discussed in Section 5.3. Using Llama-2-7b-chat, we modified Head-26 of Layer-2 for three different inputs from the AdvBench dataset. The results show that using mean ablation produces results similar to those obtained with $\epsilon$ scaling, but with some subtle differences. Specifically, ablating the Value ($W_{v}$, column 3) still has no effect on the attention weights. However, modifying the Query ($W_{q}$) and Key ($W_{k}$) no longer yields equivalent results, and the attention weights do not converge to $\frac{1}{N}$ as expected.
@@ Appendix D Attributing Safety Parameters Should Not Set the System Prompts
helpfully as possible, while being safe. Your answers should not include any
harmful, unethical, racist, sexist, toxic, dangerous, or illegal content.
@@
nature. If a question does not make any sense, or is not factually coherent,
explain why instead of answering something not correct. If you don’t know
[…]
@@
2. Avoid using harmful, unethical, racist, sexist, toxic, or illegal content.
3. Explain why instead of answering a question that does not make any sense.

### v2

@@ 4.3 How Does Safety Heads Affect Safety?
Our Method Localizes Safety Parameters at a Finer Granularity. Previous research on interpretability (Zou et al., 2023a; Xu et al., 2024c), such as ActSVD (Wei et al., 2024b), Generation-Time Activation Contrasting (GTAC) & Dynamic Activation Patching (DAP) (Chen et al., 2024) and Layer-Specific Pruning (LSP) (Zhao et al., 2024b), has identified safety-related parameters or representations. However, our method offers a more precise localization, as detailed in Table 1. We significantly narrow down the focus from parameters constituting over 5% to mere 0.018% (three heads), improving attribution precision under similar ASR by three orders of magnitude compared to existed methods.
@@ 5.1 Different Impact between Attention Weight and Attention Output
Table 3: The impact of the number of ablated safety attention heads on ASR. Upper. Results of attributing safety heads at the dataset level using generalized Ships; Bottom. Results of attributing specific harmful queries using Ships.
@@ 5.2 Pre-training is Important For LLM Safety
To explore the association between safety within attention heads and the pre-training phase, we conduct an experiment where we load the attention parameters from the base model while keeping the other parameters from the aligned model. We evaluate the safety of this ‘concatenated’ model and discover that it retains safety capability close to that of the aligned model, as shown in Figure 6(a). This observation further supports the notion that the safety effect of the attention mechanism is primarily derived from the pre-training phase. Specifically, reverting parameters to the pre-alignment state does not significantly diminish safety capability, whereas ablating a safety head does.
@@ 5.3 Helpful-Harmless Trade-off
The neurons in LLMs exhibit superposition and polysemanticity (Templeton, 2024), meaning they are often activated by multiple forms of knowledge and capabilities. Therefore, we evaluate the impact of safety heads ablation on helpfulness. We use lm-eval (Gao et al., 2024) to assess model performance after ablating safety heads of Llama-2-7b-chat on zero-shot tasks, including BoolQ (Clark et al., 2019a), RTE (Wang, 2018), WinoGrande (Sakaguchi et al., 2021), ARC Challenge (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018). As shown in Figure 6(b), we find that safety head ablation significantly degrades the safety capability while causing little helpfulness compromise. Based on this, we argue that the safety head is indeed primarily responsible for safety.
@@ 6 Conclusion
7 Acknowledgements
This work was supported by Alibaba Research Intern Program.
@@ A.1 Undifferentiated Attention
Finally, the output $h_{i}$ of the attention head degenerates to the matrix $Ah_{i}$, whose elements are the reciprocals of the number of non-zero elements in each row, which holds exactly when $\epsilon=0$.
@@ A.2 Modifying the Value Matrix Reduces the Contribution
In Figure 8, we visualize the attention weights after applying the mean operation to Query ($W_{q}$), Key ($W_{k}$), and Value ($W_{v}$), as discussed in Section 5.3. Using Llama-2-7b-chat, we modified Head-26 of Layer-2 for three different inputs from the AdvBench dataset. The results show that using mean ablation produces results similar to those obtained with $\epsilon$ scaling, but with some subtle differences. Specifically, ablating the Value ($W_{v}$, column 3) still has no effect on the attention weights. However, modifying the Query ($W_{q}$) and Key ($W_{k}$) no longer yields equivalent results, and the attention weights do not converge to $A$ as expected.
@@ Appendix D Attributing Safety Parameters Should Not Set the System Prompts
helpfully as possible, while being safe. Your answers should not include
any harmful, unethical, racist, sexist, toxic, dangerous, or illegal
content.
@@
nature. If a question does not make any sense, or is not factually
coherent,  explain why instead of answering something not correct. If you
don’t  know  the answer to a question, please don’t share false infor-
@@
2. Avoid using harmful, unethical, racist, sexist, toxic, or illegal
content.
3. Explain why instead of answering a question that does not make any
sense.
@@ Appendix E Safety Course Correction Capability Compromise
Appendix F Related Works and Discussion
LLM safety interpretability is an emerging field aimed at understanding the mechanisms behind LLM behaviors, particularly their responses to harmful queries. It is significant that understanding why LLMs still respond to harmful questions based on interpretability technique, and this view is widely accepted (Zhao et al., 2024a; Bereska & Gavves, 2024; Zheng et al., 2024c). However, dissecting the inner workings of LLMs and performing meaningful attributions remains a challenge.
RepE (Zou et al., 2023a) stands as one of the early influential contributions to safety interpretability. In early 2024, the field saw further advancements, enabling deeper exploration into this area. Notably, a pioneering study analyzed GPT-2’s toxicity shifts before and after alignment (DPO), attributing toxic generations to specific neurons (Lee et al., 2024). In contrast, our work focuses on the inherent parameters of aligned models, examining the model itself rather than focusing solely on changes. Another early approach aimed to identify a safe low-rank matrix across the entire parameter space (Wei et al., 2024b) , whereas our analysis zooms in on the multi-head attention mechanism.
Drawing inspiration from works analyzing high-level safety representations (Zheng et al., 2024a), several subsequent studies (Zhao et al., 2024b; Leong et al., 2024; Xu et al., 2024c; Zhou et al., 2024) have explored safety across different layers in LLMs. Additionally, other works (Arditi et al., 2024; Templeton, 2024) have approached safety from the residual stream perspective.
Neverthless, these works did not fully address the role of multi-head attention in model safety, which is the focus of our study. Although some mentioned attention heads, their ablation methods were insufficient for uncovering the underlying issues. Our novel ablation method provides a more effective approach for identifying safe attention heads, which constitutes a significant contribution of this paper.